Exotherimic Reacting Euler Equations

David Wagner University of Houston

We present technical improvements in the global existence of entropy solutions for exothermically reacting Euler equations. In particular we recognize that the lower order reation term causes the solution to drift in U-space, and this drift imposes conditions on the initial data. In addition we derive a relationship between ϵ or γ −1 and the total variation of the initial data; this relationship is more explicit than the one presented in Temple’s thesis.